# Attention layers provably solve single-location regression

### v1

@@ 6. Outline of the proof of Theorem 5
Notation.  In this section, we consider a constant inverse temperature schedule $\lambda_{t}\equiv\lambda>0$, as in Theorem 5. For this reason, it is not necessary to make explicit the dependence of ${\mathcal{R}}_{\lambda}$ and ${\mathcal{R}}_{\lambda}^{<}$ on $\lambda$, and we use the lighter notations ${\mathcal{R}}$ and ${\mathcal{R}}^{<}$ instead.
@@
Step 1: Invariant manifold & reparameterization.  We first show that the risk $\mathcal{R}(k,v)$ has a simpler expression when considered on the manifold $\mathcal{M}$.
@@
Step 2: Analysis of the stationary points.  Regarding the dynamics restricted to the invariant manifold $\mathcal{M}$, we can characterize the limit points of the PGD iterates as follows.
@@ 3. $(iii)$
Figure 6. Dynamics in $(\kappa,\nu)$ on the manifold $\mathcal{M}$. In (a), the fixed points of the dynamics are represented; the minimizers, saddle point, and maximizers are respectively depicted in yellow, blue and red. In (b), the vector field $(\kappa,\nu)\mapsto-(\partial_{\kappa}\mathcal{R}^{<}(\kappa,\nu)(1-\kappa^{2}),\partial_{\nu}\mathcal{R}^{<}(\kappa,\nu)(1-\nu^{2}))$ is displayed (the colormap corresponds to the magnitude of the vector field).
Step 3: Convergence to global minima.  The convergence of the sequence of iterates $(\kappa_{t},\nu_{t})_{t\geqslant 0}$ to a global minimum is shown in two stages. First, we show that the iterates converge to one of the five fixed points described in Proposition 10.
@@
These properties enable us to apply the Center-Stable Manifold theorem (Shub, 1987, Theorem III.7) to deduce the next proposition.
@@
7. Conclusion
This paper introduced single-location regression, a novel statistical task where the relevant information in the input sequence is supported by a single token. We analyzed the statistical properties and optimization dynamics of a natural estimator for this task, which simplifies to a basic attention layer. We hope this work encourages further research into how Transformer architectures address sparsity and long-range dependencies, while simultaneously constructing internal linear representations of their input—–an aspect with significant implications for interpretability. Beyond NLP, potential applications include problems connected to sparse sequential modeling such as anomaly detection in time series. A natural extension of our framework is when relevant information is spread across a few input tokens rather than just one, which relates to multi-head attention. Future mathematical analyses should consider extensions to general initialization schemes and stochastic dynamics. Our numerical experiments (Figures 4(a) and 5) yield encouraging results in both of these directions.
Acknowledgments
Authors thank Peter Bartlett, Linus Bleistein, Alex Damian, Spencer Frei, and Clément Mantoux for fruitful discussions and feedback. P.M. is supported by a Google PhD Fellowship. R.B. acknowledges support from the ANR and the Ministère de l’Enseignement Supérieur et de la Recherche.
Appendix
Organization of the Appendix.  The intermediate results of the proof of Theorem 5 in Section 6, as well as the other statements of the main text, are proven in Section A. Section B provides an expression of the risk ${\mathcal{R}}$ beyond the manifold ${\mathcal{M}}$ that extends the one provided in Lemma 6 on ${\mathcal{M}}$. Section C gives some useful technical lemmas. Finally, experimental details and additional results are in Section D.
Notation.  In the whole Appendix, we consider a constant inverse temperature schedule $\lambda_{t}\equiv\lambda>0$, as in Theorem 5. For this reason, it is not necessary to make explicit the dependence of ${\mathcal{R}}_{\lambda}$ and ${\mathcal{R}}_{\lambda}^{<}$ on $\lambda$, and we use the lighter notations ${\mathcal{R}}$ and ${\mathcal{R}}^{<}$ instead.
Appendix A Proofs of the main results
A.1. Proof of Lemma 6 and Theorem 1
@@
$J_{0}\sim\mathcal{U}(\{1,\ldots,L\})\quad\text{and}\quad\left\{\begin{array}[]{lll}X_{J_{0}}&\sim&{\mathcal{N}}\Big{(}\sqrt{\frac{d}{2}}k^{\star},\gamma^{2}I_{d}\Big{)}\\ X_{\ell}&\sim&{\mathcal{N}}(0,I_{d})\quad\textnormal{for}\quad\ell\neq J_{0}.\end{array}\right.$
@@
This last expression only involves the five parameters $\kappa,\nu,\theta,\eta,\rho$, which play a role either explicitly in the function or as parameters of the covariance of the random variables. This proves the first statement of Theorem 1. A computation of a closed-form formula for this expectation is given in Appendix B.
@@ A.4. Proof of Lemma 4
$\partial_{\theta}{\mathcal{R}}^{<}(\kappa,\nu,\theta,\eta,\rho)=-{\mathcal{R}}^{<}(\kappa,\nu,-\theta,-\eta,-\rho)\,.$
@@ A.9. Proof of Proposition 12
The mapping $g$ is a local diffeomorphism around $(0,0)$, whose Jacobian matrix has one eigenvalue in $(0,1)$ and one eigenvalue in $(1,\infty)$.  Consider the Taylor expansion of the first component $g(\kappa,\nu)_{1}$ of $g(\kappa,\nu)$. Since $\partial_{\kappa}\mathcal{R}^{<}(0,0)=0$, and $\mathcal{R}^{<}$ is smooth, letting $x=(\kappa,\nu)$, we have $(\partial_{\kappa}\mathcal{R}^{<}(\kappa,\nu))^{2}=O(\|x\|^{2})$. Thus,
@@
The mapping $g$ is differentiable on $[-1,1]^{2}$, and its Jacobian is not degenerate.  The mapping $g$ is clearly differentiable as a composition of differentiable function. The more delicate part is to show that its Jacobian cannot be degenerate. To show this statement, observe first that, for $x\in[-1,1]^{2}$, we may write $g(x)=x+\alpha h(x)$, where the first component of $h$ is given by
@@
The mapping $g$ is injective.  The computation above shows that $h$ is $\beta$-Lipschitz continuous with $\beta$ independent of $\alpha$ (for $\alpha$ small enough). In particular we can choose $\alpha$ such that $\alpha<1/\beta$. Now, let $x\neq y\in[-1,1]^{2}$ be such that $g(x)=g(y)$. Then
@@ A.10. Proof of Proposition 13
Appendix B Expression of the risk beyond the invariant manifold
In this appendix, we provide an expression of the risk ${\mathcal{R}}$ beyond the manifold ${\mathcal{M}}$ that extends the one provided in Lemma 6. This result is not needed to prove Theorem 5, and its proof is more involved that the one of Lemma 6. However, we provide it since it might be relevant to follow-up works that would study the dynamics if not initialized on the invariant manifold ${\mathcal{M}}$. It is also useful for the numerical simulations (see Appendix D).
@@ Appendix B Expression of the risk beyond the invariant manifold
A first decomposition.  We start back from the expression (15) obtained for the risk. By expanding in $\xi$, then expanding the square, we obtain
@@
$\quad+\underbrace{\sum_{\ell\neq j\geqslant 2}^{L}\mathbb{E}\Big{(}X_{\ell}^{\top}v\,\mathrm{erf}(\lambda X_{\ell}^{\top}k)X_{j}^{\top}v\,\mathrm{erf}(\lambda X_{j}^{\top}k)\Big{)}}_{=:\color[rgb]{0.68,0.36,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.68,0.36,1}\pgfsys@color@cmyk@stroke{0.32}{0.64}{0}{0}\pgfsys@color@cmyk@fill{0.32}{0.64}{0}{0}R_{3}}$
$\quad\underbrace{-2\sum_{\ell=2}^{L}\mathbb{E}\Big{(}\big{(}X_{1}^{\top}v^{\star}-X_{1}^{\top}v\,\mathrm{erf}(\lambda X_{1}^{\top}k)\big{)}X_{\ell}^{\top}v\,\mathrm{erf}(\lambda X_{\ell}^{\top}k)\Big{)}}_{=:\color[rgb]{1,0.47,0.62}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.47,0.62}\pgfsys@color@cmyk@stroke{0}{0.53}{0.38}{0}\pgfsys@color@cmyk@fill{0}{0.53}{0.38}{0}R_{4}}+\,\varepsilon^{2}\,.$
Computation of $R_{1}$.  By expanding the square,
@@
$\mathbb{E}\Big{(}\big{(}X_{1}^{\top}v^{\star}\big{)}^{2}\Big{)}=\Big{(}\mathbb{E}\big{(}X_{1}^{\top}v^{\star}\big{)}\Big{)}^{2}+\text{Var}\big{(}X_{1}^{\top}v^{\star}\big{)}=(\sqrt{\frac{d}{2}}(k^{\star})^{\top}v^{\star})^{2}+\gamma^{2}={\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\gamma^{2}}\,.$
@@
$={\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}\gamma^{2}\nu\mathrm{erf}\left(\frac{\lambda\sqrt{d/2}\kappa}{\sqrt{1+2\lambda^{2}\gamma^{2}}}\right)+\sqrt{\frac{d}{2}}\frac{\lambda\gamma^{2}\eta\theta}{\sqrt{1+2\lambda^{2}\gamma^{2}}}\mathrm{erf}^{\prime}\left(\frac{\lambda\sqrt{d/2}\kappa}{\sqrt{1+2\lambda^{2}\gamma^{2}}}\right)}$
$\qquad{\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}+\frac{\lambda^{2}\gamma^{4}\eta\rho}{1+2\gamma^{2}\lambda^{2}}\mathrm{erf}^{\prime\prime}\left(\frac{\lambda\sqrt{d/2}\kappa}{\sqrt{1+2\lambda^{2}\gamma^{2}}}\right)}$
@@
$={\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}(\frac{d}{2}\theta^{2}+\gamma^{2})\zeta\Big{(}\lambda\sqrt{\frac{d}{2}}\,\kappa,\lambda^{2}\gamma^{2}\Big{)}}$
${\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}+\sqrt{\frac{d}{2}}\left(\theta\rho-\frac{\lambda^{2}\gamma^{2}\rho^{2}\kappa}{1+2\lambda^{2}\gamma^{2}}\right)\frac{4\lambda\gamma^{2}}{\sqrt{1+2\lambda^{2}\gamma^{2}}}\mathrm{erf}\bigg{(}\frac{\lambda\sqrt{d/2}\,\kappa}{\sqrt{(1+4\lambda^{2}\gamma^{2})(1+2\lambda^{2}\gamma^{2})}}\bigg{)}\mathrm{erf}^{\prime}\bigg{(}\frac{\lambda\sqrt{d/2}\,\kappa}{\sqrt{1+2\lambda^{2}\gamma^{2}}}\bigg{)}}$
${\color[rgb]{0,0.5,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.5,1}\pgfsys@color@cmyk@stroke{1}{0.50}{0}{0}\pgfsys@color@cmyk@fill{1}{0.50}{0}{0}+\frac{2\lambda^{2}\gamma^{4}\rho^{2}}{1+2\lambda^{2}\gamma^{2}}\left(\frac{2}{\sqrt{\pi}\sqrt{1+4\lambda^{2}\gamma^{2}}}\mathrm{erf}^{\prime}\bigg{(}-\frac{\lambda\sqrt{d}\kappa}{\sqrt{1+4\lambda^{2}\gamma^{2}}}\bigg{)}\right)}$
@@
$={\color[rgb]{0,0.75,0.16}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.75,0.16}\pgfsys@color@cmyk@stroke{0.92}{0}{0.59}{0.25}\pgfsys@color@cmyk@fill{0.92}{0}{0.59}{0.25}(L-1)\left[\zeta(0,\lambda^{2})+\frac{8\lambda^{2}}{\pi\sqrt{4\lambda^{2}+1}(1+2\lambda^{2})}\rho^{2}\right]}\,.$
Computation of $R_{3}$.  Regarding the cross-product terms, by independence of the $(X_{\ell})$’s and Stein’s lemma, one gets
@@
${\color[rgb]{0.68,0.36,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.68,0.36,1}\pgfsys@color@cmyk@stroke{0.32}{0.64}{0}{0}\pgfsys@color@cmyk@fill{0.32}{0.64}{0}{0}R_{3}=\frac{4\lambda^{2}}{(1+2\lambda^{2})\pi}(L-1)(L-2)\rho^{2}}.$
Computation of $R_{4}$.  We have, again by independence and Stein’s lemma,
@@
${\color[rgb]{1,0.47,0.62}\definecolor[named]{pgfstrokecolor}{rgb}{1,0.47,0.62}\pgfsys@color@cmyk@stroke{0}{0.53}{0.38}{0}\pgfsys@color@cmyk@fill{0}{0.53}{0.38}{0}R_{4}=\frac{4\lambda(L-1)\rho}{\sqrt{(1+2\lambda^{2})\pi}}\left(\sqrt{\frac{d}{2}}\theta\,\mathrm{erf}\Big{(}\frac{\lambda\sqrt{d/2}\kappa}{\sqrt{1+2\lambda^{2}\gamma^{2}}}\Big{)}+\frac{\lambda\gamma^{2}\rho}{\sqrt{1+2\gamma^{2}\lambda^{2}}}\,\mathrm{erf}^{\prime}\Big{(}\frac{\lambda\sqrt{d/2}\kappa}{\sqrt{1+2\lambda^{2}\gamma^{2}}}\Big{)}\right)}.$
All in all.  Putting everything together, we obtain

### v2

@@ Acknowledgments
In the whole Appendix, we consider a constant inverse temperature schedule $\lambda_{t}\equiv\lambda>0$, as in Theorem 5. For this reason, it is not necessary to make explicit the dependence of ${\mathcal{R}}_{\lambda}$ and ${\mathcal{R}}_{\lambda}^{<}$ on $\lambda$, and we use the lighter notations ${\mathcal{R}}$ and ${\mathcal{R}}^{<}$ instead.
Appendix A Outline of the proof of Theorem 5
This section outlines the essential steps for the proof of Theorem 5. For clarity, the proofs are to be found in Appendix B, except the proof of Proposition 10.
Step 1: Invariant manifold & reparameterization.
@@ Appendix A Outline of the proof of Theorem 5
We first show that the risk $\mathcal{R}(k,v)$ has a simpler expression when considered on the manifold $\mathcal{M}$.
@@
Step 2: Analysis of the stationary points.
Regarding the dynamics restricted to the invariant manifold $\mathcal{M}$, we can characterize the limit points of the PGD iterates as follows.
@@ 3. $(iii)$
Figure 6: Dynamics in $(\kappa,\nu)$ on the manifold $\mathcal{M}$. In (a), the fixed points of the dynamics are represented; the minimizers, saddle point, and maximizers are respectively depicted in yellow, blue and red. In (b), the vector field $(\kappa,\nu)\mapsto-(\partial_{\kappa}\mathcal{R}^{<}(\kappa,\nu)(1-\kappa^{2}),\partial_{\nu}\mathcal{R}^{<}(\kappa,\nu)(1-\nu^{2}))$ is displayed (the colormap corresponds to the magnitude of the vector field).
Step 3: Convergence to global minima.
The convergence of the sequence of iterates $(\kappa_{t},\nu_{t})_{t\geqslant 0}$ to a global minimum is shown in two stages. First, we show that the iterates converge to one of the five fixed points described in Proposition 10.
@@
These properties enable us to apply the Center-Stable Manifold theorem (Shub, 1987, Theorem III.7), a tool from dynamical systems theory, to deduce the next proposition.
@@
The use of the Center-Stable Manifold theorem is crucial to our proof. Unfortunately, this tool does not provide quantitative rates of convergence. Obtaining a rate is a challenging task as it would require quantifying the distance of the iterates to the saddle points of the risk (the dynamics is indeed slower near saddle points), which in turn requires other tools of analysis and potentially additional assumptions.
Appendix B Proofs of the main results
B.1 Proof of Lemma 6 and Theorem 1
@@
$J_{0}\in\mathcal{P}(\{1,\ldots,L\})\quad\text{and}\quad\left\{\begin{array}[]{lll}X_{J_{0}}&\sim&{\mathcal{N}}\Big{(}\sqrt{\frac{d}{2}}k^{\star},\gamma^{2}I_{d}\Big{)}\\ X_{\ell}&\sim&{\mathcal{N}}(0,I_{d})\quad\textnormal{for}\quad\ell\neq J_{0}.\end{array}\right.$
@@
This last expression only involves the five parameters $\kappa,\nu,\theta,\eta,\rho$, which play a role either explicitly in the function or as parameters of the covariance of the random variables. This proves the first statement of Theorem 1. A computation of a closed-form formula for this expectation is given in Appendix C.
@@ B.4 Proof of Lemma 4
$\partial_{\theta}{\mathcal{R}}^{<}(\kappa,\nu,\theta,\eta,\rho)=-\partial_{\theta}{\mathcal{R}}^{<}(\kappa,\nu,-\theta,-\eta,-\rho)\,.$
@@ B.9 Proof of Proposition 12
The mapping $g$ is a local diffeomorphism around $(0,0)$, whose Jacobian matrix has one eigenvalue in $(0,1)$ and one eigenvalue in $(1,\infty)$.
Consider the Taylor expansion of the first component $g(\kappa,\nu)_{1}$ of $g(\kappa,\nu)$. Since $\partial_{\kappa}\mathcal{R}^{<}(0,0)=0$, and $\mathcal{R}^{<}$ is smooth, letting $x=(\kappa,\nu)$, we have $(\partial_{\kappa}\mathcal{R}^{<}(\kappa,\nu))^{2}=O(\|x\|^{2})$. Thus,
@@
The mapping $g$ is differentiable on $[-1,1]^{2}$, and its Jacobian is not degenerate.
The mapping $g$ is clearly differentiable as a composition of differentiable function. The more delicate part is to show that its Jacobian cannot be degenerate. To show this statement, observe first that, for $x\in[-1,1]^{2}$, we may write $g(x)=x+\alpha h(x)$, where the first component of $h$ is given by
@@
The mapping $g$ is injective.
The computation above shows that $h$ is $\beta$-Lipschitz continuous with $\beta$ independent of $\alpha$ (for $\alpha$ small enough). In particular we can choose $\alpha$ such that $\alpha<1/\beta$. Now, let $x\neq y\in[-1,1]^{2}$ be such that $g(x)=g(y)$. Then
@@ B.10 Proof of Proposition 13
Appendix C Expression of the risk beyond the invariant manifold
In this appendix, we provide an expression of the risk ${\mathcal{R}}$ beyond the manifold ${\mathcal{M}}$ that extends the one provided in Lemma 6. This result is not needed to prove Theorem 5, and its proof is more involved that the one of Lemma 6. However, we provide it since it might be relevant to follow-up works that would study the dynamics if not initialized on the invariant manifold ${\mathcal{M}}$. It is also useful for the numerical simulations (see Appendix E).
@@ Appendix C Expression of the risk beyond the invariant manifold
A first decomposition.
We start back from the expression (15) obtained for the risk. By expanding in $\xi$, then expanding the square, we obtain
@@
$\quad+\underbrace{\sum_{\ell\neq j\geqslant 2}^{L}\mathbb{E}\Big{(}X_{\ell}^{\top}v\,\mathrm{erf}(\lambda X_{\ell}^{\top}k)X_{j}^{\top}v\,\mathrm{erf}(\lambda X_{j}^{\top}k)\Big{)}}_{=:\color[rgb]{.73,.332,.828}\definecolor[named]{pgfstrokecolor}{rgb}{.73,.332,.828}R_{3}}$
$\quad\underbrace{-2\sum_{\ell=2}^{L}\mathbb{E}\Big{(}\big{(}X_{1}^{\top}v^{\star}-X_{1}^{\top}v\,\mathrm{erf}(\lambda X_{1}^{\top}k)\big{)}X_{\ell}^{\top}v\,\mathrm{erf}(\lambda X_{\ell}^{\top}k)\Big{)}}_{=:\color[rgb]{.98,.5,.448}\definecolor[named]{pgfstrokecolor}{rgb}{.98,.5,.448}R_{4}}+\,\varepsilon^{2}\,.$
Computation of $R_{1}$.
By expanding the square,
@@
$\mathbb{E}\Big{(}\big{(}X_{1}^{\top}v^{\star}\big{)}^{2}\Big{)}=\Big{(}\mathbb{E}\big{(}X_{1}^{\top}v^{\star}\big{)}\Big{)}^{2}+\text{Var}\big{(}X_{1}^{\top}v^{\star}\big{)}=(\sqrt{\frac{d}{2}}(k^{\star})^{\top}v^{\star})^{2}+\gamma^{2}={\color[rgb]{.255,.41,.884}\definecolor[named]{pgfstrokecolor}{rgb}{.255,.41,.884}\gamma^{2}}\,.$
@@
$={\color[rgb]{.255,.41,.884}\definecolor[named]{pgfstrokecolor}{rgb}{.255,.41,.884}\gamma^{2}\nu\mathrm{erf}\left(\frac{\lambda\sqrt{d/2}\kappa}{\sqrt{1+2\lambda^{2}\gamma^{2}}}\right)+\sqrt{\frac{d}{2}}\frac{\lambda\gamma^{2}\eta\theta}{\sqrt{1+2\lambda^{2}\gamma^{2}}}\mathrm{erf}^{\prime}\left(\frac{\lambda\sqrt{d/2}\kappa}{\sqrt{1+2\lambda^{2}\gamma^{2}}}\right)}$
$\qquad{\color[rgb]{.255,.41,.884}\definecolor[named]{pgfstrokecolor}{rgb}{.255,.41,.884}+\frac{\lambda^{2}\gamma^{4}\eta\rho}{1+2\gamma^{2}\lambda^{2}}\mathrm{erf}^{\prime\prime}\left(\frac{\lambda\sqrt{d/2}\kappa}{\sqrt{1+2\lambda^{2}\gamma^{2}}}\right)}$
@@
$={\color[rgb]{.255,.41,.884}\definecolor[named]{pgfstrokecolor}{rgb}{.255,.41,.884}(\frac{d}{2}\theta^{2}+\gamma^{2})\zeta\Big{(}\lambda\sqrt{\frac{d}{2}}\,\kappa,\lambda^{2}\gamma^{2}\Big{)}}$
${\color[rgb]{.255,.41,.884}\definecolor[named]{pgfstrokecolor}{rgb}{.255,.41,.884}+\sqrt{\frac{d}{2}}\left(\theta\rho-\frac{\lambda^{2}\gamma^{2}\rho^{2}\kappa}{1+2\lambda^{2}\gamma^{2}}\right)\frac{4\lambda\gamma^{2}}{\sqrt{1+2\lambda^{2}\gamma^{2}}}\mathrm{erf}\bigg{(}\frac{\lambda\sqrt{d/2}\,\kappa}{\sqrt{(1+4\lambda^{2}\gamma^{2})(1+2\lambda^{2}\gamma^{2})}}\bigg{)}\mathrm{erf}^{\prime}\bigg{(}\frac{\lambda\sqrt{d/2}\,\kappa}{\sqrt{1+2\lambda^{2}\gamma^{2}}}\bigg{)}}$
${\color[rgb]{.255,.41,.884}\definecolor[named]{pgfstrokecolor}{rgb}{.255,.41,.884}+\frac{2\lambda^{2}\gamma^{4}\rho^{2}}{1+2\lambda^{2}\gamma^{2}}\left(\frac{2}{\sqrt{\pi}\sqrt{1+4\lambda^{2}\gamma^{2}}}\mathrm{erf}^{\prime}\bigg{(}-\frac{\lambda\sqrt{d}\kappa}{\sqrt{1+4\lambda^{2}\gamma^{2}}}\bigg{)}\right)}$
@@
$={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}(L-1)\left[\zeta(0,\lambda^{2})+\frac{8\lambda^{2}}{\pi\sqrt{4\lambda^{2}+1}(1+2\lambda^{2})}\rho^{2}\right]}\,.$
Computation of $R_{3}$.
Regarding the cross-product terms, by independence of the $(X_{\ell})$’s and Stein’s lemma, one gets
@@
${\color[rgb]{.73,.332,.828}\definecolor[named]{pgfstrokecolor}{rgb}{.73,.332,.828}R_{3}=\frac{4\lambda^{2}}{(1+2\lambda^{2})\pi}(L-1)(L-2)\rho^{2}}.$
Computation of $R_{4}$.
We have, again by independence and Stein’s lemma,
@@
${\color[rgb]{.98,.5,.448}\definecolor[named]{pgfstrokecolor}{rgb}{.98,.5,.448}R_{4}=\frac{4\lambda(L-1)\rho}{\sqrt{(1+2\lambda^{2})\pi}}\left(\sqrt{\frac{d}{2}}\theta\,\mathrm{erf}\Big{(}\frac{\lambda\sqrt{d/2}\kappa}{\sqrt{1+2\lambda^{2}\gamma^{2}}}\Big{)}+\frac{\lambda\gamma^{2}\rho}{\sqrt{1+2\gamma^{2}\lambda^{2}}}\,\mathrm{erf}^{\prime}\Big{(}\frac{\lambda\sqrt{d/2}\kappa}{\sqrt{1+2\lambda^{2}\gamma^{2}}}\Big{)}\right)}.$
All in all.
Putting everything together, we obtain